The total electric flux from a cubical box of side 25.0 cm is 1.85×103 N⋅m2/C .

Part A

What charge is enclosed by the box?

Express your answer to three significant figures and include the appropriate units.
Answer #1

using guass's law,

flux is -

\phi = \intE.dA = q / e0

q / e0 = 1.85*10^3

q = 1.85*10^3 * 8.85*10^(-12)

q = 16.37 nC

answer